Rockness Recycling refurbishes rundown business students. The process uses a moving belt, which carries each student through the five steps of the process in sequence. The five steps are as follows:

STEP DESCRIPTION TIME REQUIRED
PER STUDENT
1 Unpack and place on belt 1.6 minutes
2 Strip off bad habits 1.5 minutes
3 Scrub and clean mind 0.8 minute
4 Insert modern methods 1.2 minutes
5 Polish and pack 0.6 minute

One faculty member is assigned to each of these steps. Faculty members work a 40-hour week and rotate jobs each week. Mr. Rockness has been working on a contract from General Eclectic, which requires delivery of 1,550 refurbished students per week. A representative of the human resources department has just called complaining that the company hasn’t been receiving the agreed-upon number of students. A check of finished goods inventory by Mr. Rockness reveals that there is no stock left.

a. Calculate the maximum output for each step.
Step Maximum output
(Students per week)
1        
2        
3        
4        
5        
b. Which step is the bottleneck?
step1
step2
step3
step4
step5
Answer #1

a.

Ans -

Max Capacity = 40 hrs x 60 mins = 2400 minutes

Step 1 - minute unpack and place on belt

Time required per student = 1.6 min

Max. Output = 2400/1.6 = 1500 students per week

Step 2 - strip off bad habits

Time required per student = 1.5 min

Max. Output = 2400/1.5 = 1600 students per week

Step 3 - scrub and clean mind

Time required per student = 0.8 min

Max. Output = 2400/0.8 = 3000 students per week

Step 4 - insert modern methods

Time required per student = 1.2 min

Max. Output = 2400/1.2 = 2000 students per week

Step 5 - polish and pack

Time required per student = 0.6 min

Max. Output = 2400/0.6 = 4000 students per week

b.

Ans -

Bottleneck: Least capacity in process = STEP 1

Process cannot create > what bottleneck is, so 1500 is max capacity for process as a whole.